Starting from a rest, a boat increases its speed to 5m/s with constant acceleration.

A) WHAT IS THE boat's average speed?
B) if it takes the boat 4 seconds to reach its speed how far has it traveled?
1
Expert's answer
2019-07-12T09:06:41-0400

A) We can find the boat's average speed from the formula:


"v_{avg} = \dfrac{1}{2}(v_i + v_f) = \dfrac{1}{2}(0 \dfrac{m}{s} + 5 \dfrac{m}{s}) = 2.5 \dfrac{m}{s}."

B) We can find the distance traveled by the boat from the formula:


"d = v_{avg}t = 2.5 \dfrac{m}{s} \cdot 4.0 s = 10 m."

Answer:

A) "v_{avg} = 2.5 \dfrac{m}{s}."

B) "d = 10 m."